• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D. Parno

    Log entry time 10:28:17 on November24,2008

    Entry number 250539

    keyword=Shift in BPM x

    There is a distinct shift of a little over -0.2 mm in the BPM-A measurement of the beam x position. (There is no shift in BPM-B x, and a smaller shift in the target BPM x). The Hall A tools screen displays a shift of ~0.15 mm from the desired value.

    Fig. 1 is from Run 3864, which finished at 10:00 this morning. The earliest run I could find displaying this shift was Run 3832 from Sunday morning's owl shift (Fig. 2). Prior to this run the beam position was a bit unstable.

    The trigger rates look OK.


    FIGURE 1

    FIGURE 2